how would you do this please show me {4a+2u+i=2
{5a+3u+2i=17
{ a-5u =3

Answer by kapilsinghi(68)

4a+2u+i =2 ------------1
5a+3u+2i=17 -----------2
a-5u =3 ------------3
Adding 1,2 and 3
10a + 3i = 22
10a = 22 -3i
a = 22/10 - 3i/10
according to equation 3
a-5u =3
(22-3i)/10 - 5u = 3
22-3i - 50u = 30
50u = -8 -3i
u = (-8 -3i)/50

kapilsinghi123@gmail.com